OpenStudy (anonymous):

Help??? Unit Rates>Jake is traveling 13 mph in his boat. After 3 hours, how far will he have traveled?

OpenStudy (mahoganie.carson):

What do you Think?

OpenStudy (anonymous):

39

OpenStudy (mahoganie.carson):

No

OpenStudy (mahoganie.carson):

13*5 =

OpenStudy (anonymous):

NO its 13*3+39

OpenStudy (anonymous):

In one hour he traveled 13 miles so how many will he travel in 3 hours >>>> do 3 times 13 = 39

OpenStudy (anonymous):

yes that is what i ment

OpenStudy (mahoganie.carson):

13 miles * 5 hours = 65 miles per 5 hours

OpenStudy (anonymous):

it said 3 hours

OpenStudy (anonymous):

there is no 5 in the promblem

OpenStudy (mahoganie.carson):

Oh Thanks @Nizux

OpenStudy (anonymous):

no problem

OpenStudy (mahoganie.carson):

So it would be three hours

OpenStudy (anonymous):

so @Middleschool123 u were correct the answer is 39

OpenStudy (anonymous):

\[3\times13=39 \]

OpenStudy (mahoganie.carson):

So it would be 39 miles per 3 hours

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

Good job :)

OpenStudy (mahoganie.carson):

Thanks, You Too :)

OpenStudy (mahoganie.carson):

Can you close this @middleschool123

OpenStudy (anonymous):

rate x time = distance. or r t = d t = d/r

OpenStudy (anonymous):

r = d/t

OpenStudy (mahoganie.carson):

We help him/her with the problem she/he did not close it @robtobey

OpenStudy (anonymous):

The above equations will allow one to solve most speed/time/distance problems.

OpenStudy (anonymous):

Oh